Calculate the pH of a solution made by dissolving 193 g of sodium oxalate in enough water to get a final volume of 785 mL. 


1
Expert's answer
2021-03-29T06:01:01-0400

Molar Mass of sodium Oxalate = 134

193/134 = 1.44 moles

= Concentration = 1.44/0.785

= 1.83478 mol/L

POH = -log(1.83478)

= -0.264

POH +PH = 14

PH = 14+0.264

= 14.264

= 14.3